OpenStudy (anonymous):

If h(x) = 2x + 3, find the value of x if h(x) = 9. 21 3 6 9

OpenStudy (anonymous):

You make the question h(9)=2*9+3

OpenStudy (anonymous):

IS IT 9 ? @PaulK

OpenStudy (anonymous):

no it is 9*2 + 3 = 18+3 = ?

OpenStudy (anonymous):

21

OpenStudy (anonymous):

correct :)